what will a deposit of 4500 if left in the bank for 5 years at 8% compounded quarterly?
Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

A=4500(1+0.08/4)^4*5
A=4500(1.02)^20
A=4500*1.48595
A=$6686.78

Answer by stanbon(75887)

what will a deposit of 4500 if left in the bank for 5 years at 8% compounded quarterly?
-------------------------
A(t) = P(1+(r/n))^(nt)
-----
A(5) = 4500(1+(0.08/4))^(4*5)
-----
A(5) = 4500(1.02)^20 = $6686.73
=======================================
Cheers,
Stan H.